The following questions refer to an aqueous 0.085 M sodium acetate solution.

1. What is the equilibrium concentration of [Na+]?

2. What is the equilibrium concentration of the acetate ions?

3. What is the pH of the solution?
Answer #1

1)

equilibrium concentration of [Na+]   = 0.085 M

2) CH3COO- + H2O ---------------------> CH3COOH + OH-

      0.085                                                      0                 0

          - x                                                    + x               +x  

0.085 -x                                                      x                  x

Kb = x^2 / 0.085-x = 5.56 x 10^-10

x^2 + 5.56 x 10^-10 x - 4.72 x 10^-11 = 0

x = 6.87 x 10^-6

[CH3COO-]   = 6.87 x 10^-6 M

3)

pOH = -log [OH-]

         = -log (6.87 x 10^-6)

        = 5.16

pH + pOH = 14

pH = 8.84